SOD | Year | Day of Year |
Time of Day |
Data Impact | System | Description |
---|---|---|---|---|---|---|
71191402 | 2025 | 089 | 07:00 | 3 | 8.01 | The ND filter correction and the Calibration range based on 2021 engineering measurements and the NGS 2019 local survey were updated in the onsite data reduction. The net effect of this change will reduce our observed range bias by 11.1 mm |
71191402 | 2024 | 064 | 21:40 | 3 | 6.01 | Re-seated internal and external BNC cable connectors on ETM |
71191402 | 2023 | 285 | 23:00 | 2 | 9.02 | Installed GPS antenna splitter for CNS and Rubidium XLDC |
71191402 | 2023 | 221 | 20:00 | 3 | 6.01 | Replaced ETM power supply |
71191402 | 2023 | 202 | 23:15 | 1 | 5.01 | Replaced dead start diode batteries |
71191402 | 2023 | 089 | 01:40 | 1 | 12.01 | Replacement MET4, sn 149703 installed, sn 109197 removed. |
71191402 | 2022 | 314 | 04:00 | 2 | 9.02 | Replaced Truetime XLDC. Replacement s/n 99173236 |
71191402 | 2022 | 260 | 00:00 | 2 | 6.01 | Operations change to 2900V for all satellites and calibrations |
71191402 | 2022 | 231 | 21:17 | 0 | 99 | Install new Dell 7920 replacement processor computer with CENTOS7 |
71191402 | 2022 | 029 | 18:00 | 1 | 12.01 | Replacement MET4, sn 109197 installed. Sn 117660 removed from service. |
71191402 | 2021 | 161 | 02:45 | 1 | 5.02 | Frequency synthesizer s/n 824432 failure. Replaced with FS s/n 2095. |
71191402 | 2021 | 147 | 06:05 | 1 | 5.02 | Laser frequency synthesizer s/n 824432 installed |
71191402 | 2021 | 002 | 20:15 | 1 | 5.02 | Laser frequency synthesizer electrically isolated and grounded. |
71191402 | 2020 | 312 | 00:31 | 1 | 6.01 | Install NASA Engineering Change 3273C. Improved grounding for the Event Timer Module. |
71191402 | 2020 | 241 | 00:13 | 1 | 5.02 | Install NASA Engineering Change 3395C. Start diode cable upgrade. Noted about a +15ns change in system delay. |
71191402 | 2020 | 219 | 01:30 | 1 | 5.02 | Install replacement frequency synthesizer. s/n:2095 |
71191402 | 2020 | 185 | 19:30 | 0 | 5.02 | Remove 1 ND from laser transmit insert. Thickness of glass removed is 0.001186 meters. GEODETICS.dat not changed. |
71191402 | 2020 | 154 | 22:00 | 0 | 99 | Install replacement data processing computer. |
71191402 | 2019 | 313 | 01:10 | 1 | 6.01 | Discriminator Cable Change. Move Cable 215 from Transmit Non-Delayed Channel-A to Receive Non-Delayed Channel-C. This enabled Receive Energy Recording |
71191402 | 2019 | 262 | 22:30 | 1 | 9.02 | New CNS clock installed s/n:801019 |
71191402 | 2019 | 236 | 19:06 | 1 | 12.01 | MET-4 replaced. s/n:117660 |
71191402 | 2019 | 166 | 02:40 | 1 | 5.02 | Install replacement laser frequency synthesizer. s/n:824432 |
71191402 | 2018 | 292 | 19:50 | 0 | 99 | Install replacement controller (servo) computer. |
71191402 | 2018 | 122 | 00:00 | 1 | 5.02 | Install replacement laser frequency synthesizer. s/n:806256 |
71191402 | 2018 | 021 | 00:00 | 2 | 5.02 | Return to 5Hz laser fire rate for all targets. |
71191402 | 2017 | 300 | 18:00 | 2 | 5.02 | Go to 10Hz laser fire rate for all targets less than LAGEOS1/2 orbital height. |
71191402 | 2017 | 292 | 20:00 | 3 | 6.01 | Changed time of flight measurement from HP5370B to Cybioms ETM-02 (serial number -012) |
71191402 | 2017 | 158 | 21:00 | 0 | 5.00 | Move laser chiller from Electronics Trailer to adjacent CMU building. |
71191402 | 2016 | 338 | 02:28 | 2 | 6.01 | Install new RCC (Range Control Card). |
71191402 | 2016 | 327 | 00:40 | 1 | 9.02 | Swap CNS Clock with another unit. s/n:80120 |
71191402 | 2016 | 126 | 20:00 | 2 | 6.01 | Install SLR Engineering Change (SCNS-SLR-EC1296). RCCard Upgrade and TCG Removal. |
71191402 | 2015 | 200 | 00:00 | 1 | 12.01 | MET-4 replaced. s/n:117555. |
71191402 | 2015 | 047 | 00:00 | 1 | 4.01 | Replace Transmit/Receive rotating mirror with partially mirrored lens. |
71191402 | 2015 | 043 | 00:00 | 1 | 4.01 | Replace telescope motor and position sensors. Both axes. |
71191402 | 2014 | 191 | 19:30 | 2 | 6.01 | Swap out HP5370B with another unit. |
71191402 | 2014 | 102 | 02:50 | 2 | 9.02 | Replacement GPS XL-DC installed. |
71191402 | 2012 | 362 | 20:59 | 1 | 12.00 | MET-4 Installation |
71191402 | 2012 | 310 | 23:26 | 1 | 99 | SATTRK and MONITOR software updated on the controller computer for MET 4 support |
71191402 | 2011 | 223 | 04:05 | 2 | 6.01 | Installation of replacement HP-5370B counter |
71191402 | 2010 | 214 | 19:00 | 2 | 9.02 | Replacement GPS XL-DC installed. |
71191402 | 2010 | 184 | 00:00 | 2 | 5.02 | LASER replaced with saturable absorber upgrade. |
Latest Data
Tracking Statistics
Logs